Evaluate the ∫(1-1/x)² dx


1
Expert's answer
2021-02-16T11:28:41-0500

Solution.

"\int(1-\frac{1}{x})^2dx=""\int(1-\frac{2}{x}+\frac{1}{x^2})dx=""\int1dx-2\int\frac{1}{x}dx+\int\frac{1}{x^2}dx=""x-2\ln{|x|}+\frac{x^{-1}}{-1}+C=""x-2\ln{|x|}-\frac{1}{x}+C,"

where "C" is some constant.

Answer. "x-2\ln{|x|}-\frac{1}{x}+C."